OpenStudy (jacksonheresogoon):

bracelet is marked down 20%. The original price is $37.50. How much is the discount?

Vocaloid (vocaloid):

(20/100) * original price = your answer

OpenStudy (perfecta):

$37.50 base price, 20% decrease 37.50 * 0.2 (20%) = 7.5 37.50 - 7.5 = 30

Vocaloid (vocaloid):

it's asking for the discount amount, not the discounted price

OpenStudy (chrishogan):

It's pretty simple.

Vocaloid (vocaloid):

the discounted amount is 20% of the original price

OpenStudy (aravindg):

No direct answering. You are being warned.

Vocaloid (vocaloid):

20% of 37.5 is not 30

OpenStudy (chrishogan):

Yeah it's $7.50 then you subtract $37.50 by the new price, to get $30. It's asking how much it decreased, not the finale value.

OpenStudy (chrishogan):

"How much is the discount? "

Vocaloid (vocaloid):

the answer still isn't 30

Vocaloid (vocaloid):

30 is the final amount, not the discount

Vocaloid (vocaloid):

discount = how much money did we save?

OpenStudy (chrishogan):

It can be read both ways, you dont need to get snarky about it.

Vocaloid (vocaloid):

not trying to be snarky :) I'm just a proud member of the Society of People Who Think Words Have Definitions

OpenStudy (chrishogan):

You're still being snarky pal. ;)